OpenStudy (anonymous):

Sam's coach is twice as old as Sam. When Sam started soccer 10 years ago, his coach was 3 times his age. How old is Sam now?

OpenStudy (oaktree):

If Sam's age is x and his coach is y, then y=2x and x-10=3y.

OpenStudy (anonymous):

Not quite. I think it comes out something like y = 2x and y-10 = 3(x-10) They were both younger ten years ago.

OpenStudy (anonymous):

so would that make sam 15...?

OpenStudy (anonymous):

I think if you use my equations, you will get Sam is 20. That means his coach is forty now, and ten years ago they were ten and thirty.